OpenStudy (anonymous):

Let A be an n x n matrix. Show that if A^(k+1) = O, then I - A is nonsingular and (I - A)^-1 = I + A + A^2 + . . . + A^k

OpenStudy (anonymous):

I think I have figured it out...

OpenStudy (loser66):

show me please

OpenStudy (loser66):

I was taught that if A is an n x n matrix, then 1/(I-A) = I + A +A^2 + A^3+.... but not relate to A^(k+1) =0

OpenStudy (anonymous):

(I - A)(I + A + A^1 + A^2 + . . . + A^k)= I(I + A + A^1 + A^2 + . . . + A^k) - A(I + A + A^1 + A^2 + . . . + A^k)

OpenStudy (anonymous):

= (I + A + A^1 + A^2 + . . . + A^k) - (A + A^2 + A^3 . . . + A^(k + 1))

OpenStudy (anonymous):

= (I - A^(k + 1)) = I - O , since A^(k+1) = O

OpenStudy (anonymous):

= I Thus, I - A is nonsingular and its inverse is equal to (I + A + A^1 + A^2 + . . . + A^k)

OpenStudy (loser66):

oh yea, you are brilliant!!

OpenStudy (anonymous):

You know, I haven't done any matrix division yet. I spent like and hour and a half thinking about this and the moment I posted a question something just clicked haha.

OpenStudy (loser66):

Yes, I can see it. :) Congrat!! I wish I have 10% of your brain

OpenStudy (loser66):

Thanks for posting. I learn a lot

OpenStudy (anonymous):

Thanks man. I see you interact a lot on here. Its great to see another person genuinely interested in mathematics.

OpenStudy (loser66):

:)